Given the following consumption fiction, C= 100 + 0.6YD calculate by how much induced consumption spending will decrease if the government increases taxes by R100


1
Expert's answer
2021-05-16T17:49:12-0400

Original Consumption function, "C = 100 + 0.6 Yd"

Lets call this consumption function be "C_1 = 100 + 0.6 Yd = 100+ 0.6(Y - T1)"

Now if new Taxes levied are "T_2 = T_1 + 100"

Therefore. new consumption function is"C_2 = 100 + 0.6 Yd = 100 + 0.6 (Y - T_2)"

"C_2 = 100 + 0.6(Y - T_1 -100)"

"= 100 + 0.6(Y - T_1) - 60 = 40 + 0.6(Y - T_1)"

Therefore, change in consumption spending is "C_2 - C_1 =40 + 0.6(Y - T_1) -100 - 0.6(Y - T_1) = -60"

Hence, as government increases taxes by "R 100" the consumption spending decreases by "R 60."

Therefore, this tax increase induces consumption spending to decrease by "R 60"